An RC circuit have a 9V source and lies at 54seconds of the time constant. Then find the voltage across a capacitor after 150 3.6 seconds of charging.


1
Expert's answer
2022-06-07T17:37:29-0400
"V_C=V_s(1-e^{-t/\tau})=9(1-e^{-1503.6/54})=9\:\rm V"